# Real-world Image Dehazing with Coherence-based Label Generator and Cooperative Unfolding Network

### v2

@@ 1 Introduction
Conventional methods [20, 24] are limited by fixed feature extractors, which struggle to handle the complexities of real haze. Although existing deep learning-based methods [37, 54] demonstrate improved performance, they face two significant challenges: (1) These methods do not accurately model the complex distribution of haze, leading to color distortion (as illustrated in fig. 1 DGUN [37]). (2) Real-world settings lack sufficient paired data for network training while optimizing the network with synthesized data brings a domain gap, limiting the generalizability of the models.
@@ 2.1 Real-world Image Dehazing
Given the characteristics of real haze, RIDCP [54] and Wang et al. [52] proposed novel haze synthesis pipelines. However, relying solely on synthetic data limits models’ robustness in real-world dehazing scenarios. Recognizing the distributional disparities between synthetic and real haze, methods like CDD-GAN [5], D4 [59], Shao et al. [47], and Li et al. [30] have utilized CycleGAN [69] for dehazing. Despite this, the challenges inherent in GAN [12] training often result in artifacts. Some approaches combine synthetic and real-world data, applying unsupervised loss to supervise real-world dehazing learning [6]. However, these losses lack sufficient precision, leading to suboptimal results. Other methods leverage pseudo-labels [32, 8, 50], but the erroneous pseudo-labels cause degrade quality.
@@ 2.2 Deep Unfolding Image Restoration
Deep Unfolding Networks (DUNs) integrate model-based and learning-based approaches [22, 23] and thus offer enhanced interpretability and flexibility compared to traditional learning-based methods. Increasingly, DUNs are being utilized for various image tasks, including image super-resolution [63], compressive sensing [61, 62], hyperspectral image reconstruction [28], and image fusion [68]. DGUN [37] proposes a general form of proximal gradient descent to learn degradation. However, it fails to decouple prior knowledge, relying solely on single-path DUN to model degradation and construct mappings, posing challenges in comprehending complex degradation. Yang and Sun first introduced DUNs to the image dehazing field and proposed PDN [57]. However, PDN does not exploit the complementary information between the dehazed image and the transmission map, resulting in detail blurring. Our CORUN optimizes the atmospheric scattering model and the image scene feature through dual proximal gradient descent, thus preventing overfitting and facilitating detail restoration.
@@ 3.1 Cooperative Unfolding Network
$\mathbf{T}_{k}=\underset{\mathbf{T}}{\arg\min}\frac{1}{2}\sum_{c\in\{R,G,B\}}\left\|\mathbf{P}^{c}-\hat{\mathbf{J}}_{k-1}^{c}\cdot\mathbf{T}+\mathbf{T}-\mathbf{I}\right\|^{2}_{2}+\phi(\mathbf{T}).$ (5)
@@
$\hat{\mathbf{T}}_{k}=\sum_{c\in\{R,G,B\}}(\mathbf{I}-\hat{\mathbf{J}}_{k-1}^{c}+\frac{\lambda_{k}}{(\mathbf{I}-\hat{\mathbf{J}}_{k-1}^{c})^{\top}})^{-1}\cdot(\mathbf{I}-\mathbf{P}^{c}\frac{\lambda_{k}\mathbf{T}_{k-1}}{(\mathbf{I}-\hat{\mathbf{J}}_{k-1}^{c})^{\top}}).$ (7)
@@
Details about CPMM. T-CPMM and S-CPMM share the same structure, which is modified from MST [2] for improved mapping quality. Each CPMM block uses a 4-channel convolution to embed $\mathbf{T}$ and $\mathbf{J}$ into a 30-dimensional feature map. The distinction between T-CPMM and S-CPMM lies in their outputs: T-CPMM produces a 1-channel result to aid TGDM in predicting a scene-compliant transmission map, whereas S-CPMM generates a 3-channel RGB image. This enables S-CPMM to learn additional scene feature information, such as atmospheric light and blur, assisting SGDM in generating higher-quality dehazed results with more details. For more efficient computation, each CPMM comprises only 3 layers with $[1,1,1]$ blocks, doubling the dimensions with increasing depth.
@@ 3.2 Coherence-based Pseudo Labeling by Colabator
Figure 3: The plug-and-play Coherence-based Pseudo Labeling paradigm.
@@
where $\mathbf{P}^{R}_{\widetilde{HQ}}\in\mathbb{R}^{H\times W\times 3}$ is the result from the teacher network using the non-augmented input, and $\mathbf{P}^{R}_{HQ}\in\mathbb{R}^{H\times W\times 3}$represents the result from the student network by strong augment input and $\mathbf{T}^{R}_{\widetilde{HQ}}$, $\mathbf{T}^{R}_{HQ}$ are the corresponding transmission map. The different degrees of data augmentation lead to varying dehazing results, typically resulting in $\mathbf{P}^{R}_{\widetilde{HQ}}$ having better quality than $\mathbf{P}^{R}_{HQ}$. This approach ensures the model descends in the correct direction and helps mitigate the overfitting issues often associated with direct pseudo-label learning methods. By iterating, our teacher network generates increasingly high-quality pseudo-labels, providing more reliable supervision for model learning.
@@
$w=\varPsi(\text{norm}(\mathcal{D}(\mathbf{S}^{R}_{\widetilde{HQ}}))+\text{norm}(\mathcal{Q}(\mathbf{S}^{R}_{\widetilde{HQ}})),$ (12)
where $\varPsi$ is compose sequence to map and resize as $\mathbf{P}^{R}_{\widetilde{HQ}}$, ,$\text{norm}(\cdot)$ means normalize scores from 0 to 1, that higher score means lower haze density and better image quality.
@@ 3.3 Semi-supervised Real-world Image Dehazing
$\begin{split}L^{common}_{Rec}(\mathbf{P}_{HQ},\mathbf{P}_{GT})=&\|\mathbf{P}_{GT},\mathbf{P}_{HQ}\|_{1}+\beta_{c}\sum_{i=1}^{n}\tau_{i}\|\varphi_{i}(\mathbf{P}_{GT}),\varphi_{i}(\mathbf{P}_{HQ})\|_{1},\\ L^{contra}_{Rec}(\mathbf{P}_{LQ},\mathbf{P}_{HQ},\mathbf{P}_{GT})=&\|\mathbf{P}_{GT},\mathbf{P}_{HQ}\|_{1}+\beta_{c}\sum_{i=1}^{n}\tau_{i}\frac{\|\varphi_{i}(\mathbf{P}_{GT}),\varphi_{i}(\mathbf{P}_{HQ})\|_{1}}{\|\varphi_{i}(\mathbf{P}_{LQ}),\varphi_{i}(\mathbf{P}_{HQ})\|_{1}},\end{split}$ (15)
@@ 4.1 Experimental Setup
Implementation Details. Our framework is implemented using PyTorch [38] and trained on four NVIDIA RTX 4090 GPUs. During the pre-training phase, we train the network for 30K iterations, optimizing it with AdamW [34, 9] using momentum parameters $(\beta_{1}=0.9,\beta_{2}=0.999)$ and an initial learning rate of $2\times 10^{-4}$, gradually reduced to $1\times 10^{-6}$ with cosine annealing [33]. In Colabator, the initial learning rate is set to $5\times 10^{-5}$ with only 5K iterations. Following [54], we employ random crop and flip for synthetic data augmentation. We use DA-CLIP [35] as our haze density evaluator and MUSIQ [25] as the image quality evaluator. Our CORUN consists of 4 stages and the trade-off parameters in the loss are set to $\beta_{c},\rho_{r},\rho_{c}$ are set to $0.2,5,10^{-2},$ respectively.
@@ 4.2 Comparative Evaluation
We compare our method with 8 state-of-the-art methods: PDN [57], MBDN [10], DH [13], DAD [47], PSD [6], D4 [59], RIDCP [54], DGUN [37]. The quantitative results, presented in table 1, show that our method achieved the highest performance, outperforming the second-best method (RIDCP) by $19.0\%$. Specifically, our method improved FADE, BRISQUE, and NIMA scores by $20.4\%$, $30.8\%$, and $7.0\%$, respectively. This demonstrates that our method surpasses current state-of-the-art techniques in both dehazing capability and the quality, and aesthetics of the generated images.
@@
Datasets Metrics w/o w/o Trusted w/o Optimal
IMD weight label pool
@@ 4.3 Ablation Study
Effect of Colabator. We validate the effect of our Colabator. In table 4, we systematically removed critical components, such as iterative mean-teacher (IMD), trusted weighting, and the optimal label pool, from the model architecture. The outcomes indicate the performance deteriorates when these components are removed, highlighting their essential role in the system.
@@ 4.4 User Study and Downstream Task
Downstream Task Evaluation. The performance of high-level vision tasks, e.g. object detection and semantic segmentation, is greatly affected by image quality, with severely degraded images often leading to erroneous results [29, 51, 58]. To address this performance degradation, some methods have incorporated image restoration as a preprocessing step for high-level vision tasks. To validate the effectiveness of our approach for high-level vision, we utilized pretrained YOLOv3 [45], and tested it on the RTTS [27] dataset, and evaluated the results using the mean Average Precision (mAP) metric. As shown in table 6 and fig. 6, our method demonstrates a substantial advantage over existing methods, verifying our efficacy in facilitating high-level vision understanding.
4.5 Limitations and Future Work
@@
In fig. 7, our CORUN+ model struggles to maintain result quality and preserve texture details when dealing with severely degraded inputs, such as strong compression and extreme high-density haze. This challenge persists across existing methods and remains unresolved. We attribute this difficulty to the model’s struggle in reconstructing scenes from dense fog, where information is often severely lacking or entirely lost, affecting the reconstruction of both haze-free and low haze density areas. Moreover, the model solely focuses on defogging and lacks the capability to address other image degradations, such as image deblurring[65] and low-light image enhancement [14], limiting its ability to achieve high-quality reconstruction results from complex degraded images. To address this limitation in future research, we propose not only focusing on environmental degradation but also considering additional information about image degradation when solving real-world dehazing problems. In addition to this, we can introduce more modalities as supplements to RGB images, enhancing the model’s ability to effectively recover details.
4.6 Broader Impacts
Real-world image dehazing is a crucial task in image restoration, aimed at removing haze degradation from images captured in real-world scenarios. In computer vision, dehazing can benefit downstream tasks such as object detection [66, 40, 41, 39], image segmentation [18, 19, 55, 17], and depth estimation [4, 26], with applications ranging from autonomous driving to security monitoring. Our paper introduces a cooperative unfolding network and a plug-and-play pseudo-labeling framework, achieving state-of-the-art performance in real-world dehazing tasks. Notably, image dehazing techniques have yet to exhibit negative social impacts. Our proposed CORUN and Colabator methods also do not present any foreseeable negative societal consequences.

### v3

@@ 1 Introduction
Conventional methods [3, 4] are limited by fixed feature extractors, which struggle to handle the complexities of real haze. Although existing deep learning-based methods [5, 6, 7, 8, 9] demonstrate improved performance, they face two significant challenges: (1) These methods do not accurately model the complex distribution of haze, leading to color distortion, as illustrated in fig. 1 DGUN [10]. (2) Real-world settings lack sufficient paired data for network training while optimizing the network with synthesized data brings a domain gap, limiting the generalizability of the models.
@@ 2.1 Real-world Image Dehazing
Given the characteristics of real haze, RIDCP [7] and Wang et al. [24] proposed novel haze synthesis pipelines. However, relying solely on synthetic data limits models’ robustness in real-world dehazing scenarios. Recognizing the distributional disparities between synthetic and real haze, methods like CDD-GAN [25], D4 [26], Shao et al. [27], and Li et al. [28] have utilized CycleGAN [29] for dehazing. Despite this, the challenges inherent in GAN [30] training often result in artifacts. Some approaches combine synthetic and real-world data, applying unsupervised loss to supervise real-world dehazing learning [19]. However, these losses lack sufficient precision, leading to suboptimal results. Other methods leverage pseudo-labels [31, 32], but the erroneous pseudo-labels cause degrade quality.
@@ 2.2 Deep Unfolding Image Restoration
Deep Unfolding Networks (DUNs) integrate model-based and learning-based approaches [33, 34] and thus offer enhanced interpretability and flexibility compared to traditional learning-based methods. Increasingly, DUNs are being utilized for various image tasks, including image super-resolution [35], compressive sensing [36], and hyperspectral image reconstruction [37]. DGUN [10] proposes a general form of proximal gradient descent to learn degradation. However, it fails to decouple prior knowledge, relying solely on single-path DUN to model degradation and construct mappings, posing challenges in comprehending complex degradation. Yang and Sun first introduced DUNs to the image dehazing field and proposed PDN [11]. However, PDN does not exploit the complementary information between the dehazed image and the transmission map, resulting in detail blurring. Our CORUN optimizes the atmospheric scattering model and the image scene feature through dual proximal gradient descent, thus preventing overfitting and facilitating detail restoration.
@@ 3.1 Cooperative Unfolding Network
$\mathbf{T}_{k}=\underset{\mathbf{T}}{\arg\min}\frac{1}{2}\left\|\mathbf{P}-\hat{\mathbf{J}}_{k-1}\cdot\mathbf{T}+\mathbf{T}-\mathbf{I}\right\|^{2}_{2}+\phi(\mathbf{T}).$ (5)
@@
$\hat{\mathbf{T}}_{k}=\sum_{c\in\{R,G,B\}}(\mathbf{I}-\hat{\mathbf{J}}_{k-1}^{c}+\lambda_{k}(\mathbf{I}-\hat{\mathbf{J}}_{k-1}^{c})^{-\top})^{-1}\cdot(\mathbf{I}-\mathbf{P}^{c}\lambda_{k}\mathbf{T}_{k-1}(\mathbf{I}-\hat{\mathbf{J}}_{k-1}^{c})^{-\top}).$ (7)
@@
Details about CPMM. T-CPMM and S-CPMM share the same structure for improved mapping quality. Each CPMM block uses a 4-channel convolution to embed $\mathbf{T}$ and $\mathbf{J}$ into a 30-dimensional feature map. The distinction between T-CPMM and S-CPMM lies in their outputs: T-CPMM produces a 1-channel result to aid TGDM in predicting a scene-compliant transmission map, whereas S-CPMM generates a 3-channel RGB image. This enables S-CPMM to learn additional scene feature information, such as atmospheric light and blur, assisting SGDM in generating higher-quality dehazed results with more details. For more efficient computation, each CPMM comprises only 3 layers with $[1,1,1]$ blocks, doubling the dimensions with increasing depth.
@@ 3.2 Coherence-based Pseudo Labeling by Colabator
Figure 3: The plug-and-play Coherence-based Pseudo-label Generator paradigm.
@@
where $\mathbf{P}^{R}_{\widetilde{HQ}}\in\mathbb{R}^{H\times W\times 3}$ is the result from the teacher network using the non-augmented input, and $\mathbf{P}^{R}_{HQ}\in\mathbb{R}^{H\times W\times 3}$represents the result from the student network by strong augment input and $\mathbf{T}^{R}_{\widetilde{HQ}}$, $\mathbf{T}^{R}_{HQ}$ are the corresponding transmission map. The different degrees of data augmentation lead to varying dehazing results, typically resulting in $\mathbf{P}^{R}_{\widetilde{HQ}}$ having better quality than $\mathbf{P}^{R}_{HQ}$.
This approach ensures the model descends in the correct direction and helps mitigate the overfitting issues often associated with direct pseudo-label learning methods. By iterating, our teacher network generates increasingly high-quality pseudo-labels, providing more reliable supervision.
@@
$w=\varPsi(\text{norm}(\mathcal{D}(\mathbf{S}^{R}_{\widetilde{HQ}}))\cdot\text{norm}(\mathcal{Q}(\mathbf{S}^{R}_{\widetilde{HQ}})),$ (12)
where $\varPsi$ is compose sequence to map and resize as $\mathbf{P}^{R}_{\widetilde{HQ}}$, $\text{norm}(\cdot)$ means normalize scores from 0 to 1, that higher score means lower haze density and better image quality.
@@ 3.3 Semi-supervised Real-world Image Dehazing
$L^{common}_{Rec}(\mathbf{P}_{HQ},\mathbf{P}_{GT})=\|\mathbf{P}_{GT},\mathbf{P}_{HQ}\|_{1}+\beta_{c}\sum_{i=1}^{n}\tau_{i}\|\varphi_{i}(\mathbf{P}_{GT}),\varphi_{i}(\mathbf{P}_{HQ})\|_{1}$ (15)
$L^{contra}_{Rec}(\mathbf{P}_{LQ},\mathbf{P}_{HQ},\mathbf{P}_{GT})=\|\mathbf{P}_{GT},\mathbf{P}_{HQ}\|_{1}+\beta_{c}\sum_{i=1}^{n}\tau_{i}\frac{\|\varphi_{i}(\mathbf{P}_{GT}),\varphi_{i}(\mathbf{P}_{HQ})\|_{1}}{\|\varphi_{i}(\mathbf{P}_{LQ}),\varphi_{i}(\mathbf{P}_{HQ})\|_{1}},$ (16)
@@ 4.1 Experimental Setup
Implementation Details. Our framework is implemented using PyTorch [43] and trained on four NVIDIA RTX 4090 GPUs. During the pre-training phase, we train the network for 30K iterations, optimizing it with AdamW [44] using momentum parameters $(\beta_{1}=0.9,\beta_{2}=0.999)$ and an initial learning rate of $2\times 10^{-4}$, gradually reduced to $1\times 10^{-6}$ with cosine annealing. In Colabator, the initial learning rate is set to $5\times 10^{-5}$ with only 5K iterations. Following [7], we employ random crop and flip for synthetic data augmentation. We use DA-CLIP [45] as our haze density evaluator and MUSIQ [46] as the image quality evaluator. Our CORUN consists of 4 stages and the trade-off parameters in the loss are set to $\beta_{c},\rho_{r},\rho_{c}$ are set to $0.2,5,10^{-2},$ respectively.
@@ 4.2 Comparative Evaluation
We compare our method with 8 state-of-the-art methods: PDN [11], MBDN [14], DH [15], DAD [27], PSD [19], D4 [26], RIDCP [7], DGUN [10]. The quantitative results, presented in table 1, show that our method achieved the highest performance, outperforming the second-best method (RIDCP) by $17.0\%$. Specifically, our method improved FADE, BRISQUE, and NIMA scores by $12.7\%$, $30.8\%$, and $7.6\%$, respectively. This demonstrates that our method surpasses current state-of-the-art techniques in both dehazing capability and the quality, and aesthetics of the generated images.
@@
Datasets Metrics w/o Mean- w/o Trusted w/o Optimal
teacher weight label pool
@@ 4.3 Ablation Study
Effect of Colabator. We validate the effect of our Colabator. In table 4, we systematically removed critical components, such as mean-teacher, trusted weighting, and the optimal label pool, from the model architecture. The outcomes indicate the performance deteriorates when these components are removed, highlighting their essential role in the system.
@@ 4.4 User Study and Downstream Task
Downstream Task Evaluation. The performance of high-level vision tasks, e.g. object detection and semantic segmentation, is greatly affected by image quality, with severely degraded images often leading to erroneous results [52, 53]. To address this performance degradation, some methods have incorporated image restoration as a preprocessing step for high-level vision tasks. To validate the effectiveness of our approach for high-level vision, we utilized pretrained YOLOv3 [54], and tested it on the RTTS [40] dataset, and evaluated the results using the mean Average Precision (mAP) metric. As shown in table 6 and fig. 6, our method demonstrates a substantial advantage over existing methods, verifying our efficacy in facilitating high-level vision understanding.
5 Limitations and Future Work
@@
In fig. 7, our CORUN+ model struggles to maintain result quality and preserve texture details when dealing with severely degraded inputs, such as strong compression and extreme high-density haze. This challenge persists across existing methods and remains unresolved. We attribute this difficulty to the model’s struggle in reconstructing scenes from dense haze, where information is often severely lacking or entirely lost, affecting the reconstruction of both haze-free and low haze density areas. Moreover, the model solely focuses on dehazing and lacks the capability to address other image degradations, such as image deblurring[55] and low-light image enhancement [56, 57], limiting its ability to achieve high-quality reconstruction results from complex degraded images. To address this limitation in future research, we propose not only focusing on environmental degradation but also considering additional information about image degradation when solving real-world dehazing problems. In addition to this, we can integrate robust generative methods to improve the network’s ability to restore dense haze regions [58, 59, 60, 61, 62], synthesize haze that matches real-world distributions [63, 64, 65, 66], and introduce more modalities as supplements to RGB images, enhancing the model’s ability to effectively recover details [67].
@@ 6 Conclusions
Appendix A Appendix
Figure 8: Detailed Comparison of fig. 4. Red region reflects all past methods have had haze residues, but our method have the least in the same case. Purple region shows our method restores richer detail and truer colors. Please zoom in for a better view.
A.1 Declaration of eq. 7 and eq. 10
Having gotten eq. 5 and eq. 6, the solution of $\hat{\mathbf{T}}$ can be formulated according to the proximal gradient algorithm:
$\mathcal{T}(\hat{\mathbf{J}}_{k-1},\mathbf{T}_{k-1})=\frac{1}{2}\|\mathbf{P}-\hat{\mathbf{J}}_{k-1}\cdot\hat{\mathbf{T}}+\hat{\mathbf{T}}-\mathbf{I}\|^{2}_{2}+\frac{\lambda_{k}}{2}\|\hat{\mathbf{T}}-\mathbf{T}_{k-1}\|^{2}_{2}.$ (22)
Then we obtain the partial derivative:
$\partial_{\hat{\mathbf{T}}}\mathcal{T}(\hat{\mathbf{J}}_{k-1},\mathbf{T}_{k-1})=(\mathbf{I}-\hat{\mathbf{J}}_{k-1})^{T}(\mathbf{P}-\hat{\mathbf{J}}_{k-1}\cdot\hat{\mathbf{T}}+\hat{\mathbf{T}}-\mathbf{I})+\lambda_{k}(\hat{\mathbf{T}}-\mathbf{T}_{k-1}).$ (23)
Let the partial derivative be equal to zero, we archieve the closed-form solution for $\hat{\mathbf{T}}$ in eq. 7.
Similarly, the solution of $\hat{\mathbf{J}}$ can be formulated as
$\mathcal{J}(\hat{\mathbf{T}}_{k},\mathbf{J}_{k-1})=\frac{1}{2}\|\mathbf{P}-\hat{\mathbf{J}}\cdot\hat{\mathbf{T}}_{k}+\hat{\mathbf{T}}_{k}-\mathbf{I}\|^{2}_{2}+\frac{\mu_{k}}{2}\|\hat{\mathbf{J}}-\mathbf{J}_{k-1}\|^{2}_{2}.$ (24)
The corresponding partial derivative is
$\partial_{\hat{\mathbf{J}}}\mathcal{J}(\hat{\mathbf{T}}_{k},\mathbf{J}_{k-1})=-\hat{\mathbf{T}}_{k}^{T}(\mathbf{P}-\hat{\mathbf{J}}\cdot\hat{\mathbf{T}}_{k}+\hat{\mathbf{T}}_{k}-\mathbf{I})+\mu_{k}(\hat{\mathbf{J}}-\mathbf{J}_{k-1}).$ (25)
The closed-form solution for $\hat{\mathbf{J}}$ is presented in eq. 10 when let the partial derivative be equal to zero.
A.2 Declaration of CLIP module.
The haze density evaluator $\mathcal{D}(\cdot)$ can be formulated as:
$\mathcal{D}(\cdot)=\frac{Enc_{image}(\cdot)}{\|Enc_{image}(\cdot)\|}\cdot(\frac{Enc_{text}(\text{Text})}{\|Enc_{text}(\text{Text})\|})^{\top}$ (26)
The text we used is "hazy" from the DA-CLIP provided in the text list.
A.3 Ablation Study
Table 7: Ablation of CPMM module of our CORUN.
Modules NIMA $\uparrow$ BRISQUE $\downarrow$ FADE $\downarrow$
Hazy(Input) 4.483 36.642 2.484
w/o CPMM 4.836 38.197 1.362
w/ CPMM (CORUN+) 5.342 11.956 0.824
Table 8: Ablation of our trusted weights present as a map or value.
Methods NIMA $\uparrow$ BRISQUE $\downarrow$ FADE $\downarrow$
Only Full 5.229 13.099 0.803
Partition+Full(CORUN+) 5.342 11.956 0.824
Table 9: Effects of more Colabator components.
Modules NIMA $\uparrow$ BRISQUE $\downarrow$ FADE $\uparrow$
w/o Colabator 4.856 16.541 1.091
w/o Strong aug. 5.084 12.671 0.813
w/o DA-CLIP[45] 5.358 11.200 0.856
CORUN+ 5.342 11.956 0.824
Table 10: Ablation of mainstream datasets setting.
RIDCP[7] Pipeline - Metrics
[…]
4.845 20.779 0.765
4.991 16.478 0.840
5.342 11.956 0.824
Table 11: Ablation of our simplified ASM formula.
[…]
w/o simplify 5.203 14.469 0.817
w/ simplify(CORUN+) 5.342 11.956 0.824
Table 12: Ablations of eq.15 and eq.16 loss functions. Our strategy achieves a better result.
Loss NIMA $\uparrow$ BRISQUE $\downarrow$ FADE $\downarrow$
Eq.15 Only 5.249 13.997 1.035
Eq.16 Only 5.220 12.484 0.795
Both Loss (Ours) 5.342 11.956 0.824
Table 13: Effects of integrating our Colabator with more cutting-edge dehazing methods. The gains brought by Colabator are significant.
[…]
C2PNet[22] 4.715 34.314 2.064
C2PNet+Colabator 4.823 23.662 1.329
FFA-Net[17] 4.822 33.235 2.080
FFA-Net+Colabator 4.839 29.219 0.958
GDN[16] 5.074 33.051 2.611
GDN+Colabator 5.258 23.691 0.947
Effect of the CPMM Module in CORUN. We evaluate the impact of the Cooperative Proximal Mapping Modules (CPMM) in our CORUN architecture. As shown in table 8, incorporating CPMM leads to a significant improvement in performance across all metrics. The model with CPMM (CORUN+) outperforms the variant without CPMM, highlighting the importance of CPMM in enhancing dehazing efficiency and image quality.
Impact of Trusted Weight Representations. We investigate the effect of using trusted weights as either a map or a value. As seen in table 8, the combination of partitioned and full trusted weights (CORUN+) achieves better results compared to using only full trusted weights. This emphasizes the value of our trusted weight representation in improving the accuracy of dehazing.
Effects of Additional Colabator Components. To analyze the contribution of individual components within the Colabator framework, we performed ablation studies, with the results presented in table 10. Removing essential elements, such as strong augmentation or DA-CLIP, results in performance deterioration, confirming the importance of each Colabator component in ensuring optimal dehazing outcomes.
Ablation on Dataset Configurations. We evaluate our methods with different dataset settings. As shown in table 10, the results verify our method still achieves a leading place under the three settings compared with existing methods.
Effectiveness of the Simplified ASM Formula.We assess the impact of simplifying the Atmospheric Scattering Model (ASM) formula. The results in table 13 indicate that using the simplified ASM formula will lead to a slight decrease in the dehazing ability, but it can evidently improve the image quality of the results.
Influence of Loss Functions. We compare the effect of using different loss functions (eq. 15 and eq. 16). Table 13 shows that combining both loss functions yields better performance than using either one alone, demonstrating the advantage of this combined loss strategy in refining the dehazing process.
Integration with Other Dehazing Methods. To test the generalizability of Colabator, we integrated it with various state-of-the-art dehazing models, such as C2PNet [22], FFA-Net [17], and GDN [16]. As shown in table 13, incorporating Colabator leads to performance gains across all metrics, demonstrating its effectiveness as a plug-and-play module for improving dehazing in various architectures.
A.4 Broader Impacts
Real-world image dehazing is a crucial task in image restoration, aimed at removing haze degradation from images captured in real-world scenarios. In computer vision, dehazing can benefit downstream tasks such as object detection [68, 69, 70, 71, 72], image segmentation [73, 74, 75, 76, 77, 78, 79, 80, 81], tracking [82, 83], depth estimation [84, 85], and more vision related tasks [86, 87, 88, 89, 90, 91, 92, 93, 94], with applications ranging from autonomous driving to security monitoring. Our paper introduces a cooperative unfolding network and a plug-and-play pseudo-labeling framework, achieving state-of-the-art performance in real-world dehazing tasks. Notably, image dehazing techniques have yet to exhibit negative social impacts. Our proposed CORUN and Colabator methods also do not present any foreseeable negative societal consequences.